# Unsupervised deep learning model for fast energy layer pre-selection of delivery-efficient proton arc therapy plan optimization of nasopharyngeal carcinoma

### v1

@@ 2.4 Evaluation
A total of $54$ patients who underwent radiotherapy for NPC were retrospectively selected for evaluation by SPArcps and SPArcdl. The performance of SPArcdl across all five validation datasets ($54$ cases) were aggregated for assessment. We calculate metrics mean for quantitative evaluation and select four representative cases for qualitative evaluation. Two structures, CTV and brainstem, are selected for evaluation. Full arc gantry angle selection was employed, with an angular sampling interval of $5$ degrees due to the constraints of computing power. Our energy selection method depended on matRad supplying the initial complete set of ELs that span the entire treatment target longitudinally for every gantry angle. A spot spacing of $5$ mm was used in both the beam and lateral directions. Dose calculations were performed with a resolution of $3\times 3\times 3$ $mm^{3}$, and the voxel size of the CT images was $5$ $cm^{3}$. The prescription dose of CTV is $54$ Gy with $30$ fractions, making $1.8$ Gy per fraction.
@@ 2.4.1 Plan quality evaluation
The performance of SPArcdl in respect to plan quality of the PAT plan generated by the predicted ELs is evaluated. The mean value of D98, D2, homogeneity index (HI), conformity index (CI), mean and max dose of brainstem are calculated for comparison between SPArcdl and SPArcps.
HI quantifies how uniformly the prescribed radiation dose is distributed within the target volume. A low HI value, ideally close to 00, indicating a more homogeneous dose distribution, suggesting less dose variation within the target. Eq.(5) presents the formula of HI described in [26].
$HI=\frac{D_{2}-D_{98}}{D_{prescription}},$ (5)
@@
Figure 2: Left column are EL prediction of $4$ representative cases generated by SPArcdl. The horizontal axes represents gantry angles and vertical axes represents proton energy (Mev). Right column are CT slices with dose distribution of the corresponding cases.
@@ 2.4.2 EL switch time and model efficiency analysis
A commonly emphasized objective in existing studies is the reduction of treatment delivery time [27, 17, 12, 15, 29]. The duration of arc delivery is influenced by various machine-specific factors, such as the number and order of ELs, spot positions, and gantry angles. Among these, ELS time often contributes the most to overall delivery time, especially when the number of energy layers and energy switch-ups is not minimized [27, 17, 12]. The ELS time is evaluated and analyzed for both SPArcps and SPArcdl. The efficiency to generate EL pre-selection, or the inference time of SPArcdl, is also reported.
@@ 3 Result
3.1 Prediction
Figure 2 demonstrates the ELs of four representative cases generated by SPArcdl and the CT slices with dose distribution of the corresponding cases. While most of the predicted ELs are in the range between $80$ and $145$, many short horizontal lines can be observed, representing the ELs at these consecutive gantry angles are unchanged. The CT slices with dose distribution demonstrate a decent CTV coverage while maintaining acceptable sparing of brainstem.
Table 1: The dosimetric performance comparison of 54 cases of SPArcps and SPArcdl; D98: the minimum doses received by $98\%$ of the CTV; D2: the minimum doses received by $2\%$ of the CTV; CI: CTV conformity; HI: CTV homogeneity; Mean and Max: the mean dose and max dose for brainstem.
Structure Type SPArcps SPArcdl (Ours)
CTV D98 1.10$\pm$0.59 1.66$\pm$0.03
D2 1.98$\pm$0.13 1.83$\pm$0.01
CI 0.60$\pm$0.12 0.76$\pm$0.04
HI $\downarrow$ 0.88$\pm$0.70 0.17$\pm$0.04
BrainStem Mean $\downarrow$ 0.84$\pm$0.24 0.63$\pm$0.19
Max $\downarrow$ 1.94$\pm$0.16 1.79$\pm$0.11
@@ 3.2 Plan quality
Table 1 demonstrates the quantitative dosimetric performance comparison, including D98, D2, CI, HI, brainstem mean and max, of SPArcdl and SPArcps on nominal distribution. With regard to CTV, SPArcdl plan presents a significantly ($p<0.01$) smaller average HI compared to the SPArcps with average differences of $0.71$. The D2 values are close between the two plans with SPArcdl plan displaying a lower value, and this difference is statistically significant ($p<0.01$). These results indicate SPArcdl can provide a more homogeneous dose distribution. Furthermore, SPArcdl exhibits a higher CI value, differing by $0.16$ compared to SPArcps. The CI difference is significant ($p<0.01$), reflecting better conformity of SPArcdl to the target. In respect to brainstem, SPArcdl yields a lower mean and max dose, differing by $0.21$ and $0.15$. The mean and max difference are both statistically significant ($p<0.01$), demonstrating better sparing of SPArcdl on this critical structure.
@@
A more intuitive performance comparison of SPArcdl (blue triangle) and SPArcps (red squares) plans is demonstrated in Figure 3, in which the distribution of CI, HI, D98, D2, mean and max dose of brainstem are presented as a function of GTV using kernel density estimates and scatter plots. In plot (a), the HI values of SPArcps plan range from about $0.2$ to $2.5$, with a dense cluster around 0.2-1, while those of SPArcdl plan are lower and with a tighter cluster, mostly between $0.1$ and $0.3$ suggesting improved homogeneity of the proposed method. Also in plot (c), the D98 values of SPArcdl plan are higher, mostly between $1.5$ and $1.8$, with a dense cluster around 1.6-1.7, indicating better minimum dose coverage. The CI values of SPArcdl are higher in plot (b), mostly between $0.6$ and $0.8$, with a dense cluster around the same range, showing better conformity to the target volume. The observations and findings of Figure 3 align with the insights of Table 1. In all four plots, the relatively denser cluster generated by SPArcdl plan indicates greater consistency of its performance.
The denser clustering and reduced variance observed in SPArcdl’s results in Figure 3 demonstrate its superior capacity to generate consistent outcomes compared to SPArcps in EL pre-selection.
Figure 3: Comparison of various metrics performance of $54$ cases between SPArcdl and SPArcps plans. Each diagram shows the relationship between GTV volume and the respective metric, displaying kernel density estimate (KDE) plots and scatter points for both methods.
A comparable plan in a previous study [15] reported low dose conformity index with respect to small lung tumor sizes. Although the primary focus of this study is NPC, we are interested in checking if we have the similar finding. As demonstrated in Figure 3 (b), the CI exhibits modest decline as the GTV volume reduces from approximately $100$ to $20$ cc. In short, SPArcdl demonstrates a slightly lower CI for relatively small NPC GTV volume.
3.3 EL switch time
@@
The performance of the model is evaluated in terms of mean energy layer switch (ELS) time in Table 2. unchanged ELS refers to the scenario in which the EL remains unchanged when the beam angle transitions to the next position. In such cases, the ELS time is effectively zero, which is shorter than that required for both ELS upward or downward transitions. Table 2 presents the mean ELS associated with ascended, descended and unchanged transition of SPArcdl and SPArcps. The mean unchanged ELS of SPArcdl is $52.6$, which is dramatically higher than that of SPArcps or $0.3$. Compared to SPArcps, the number of descended ELS of SPArcdl is almost entirely shifted to the unchanged ELS. While the mean ascended ELS of SPArcdl is slightly higher than that of SPArcps, the notable advantage of SPArcdl lies in its higher frequency of unchanged ELS. This advantage not only offsets the the cost of slightly higher EL in ascended transitions but also contributes to energy switch time of $51.9$ seconds, an $38.4\%$ reduction from SPArcps. In addition, Figure 2 demonstrates a few cases of EL generated by SPArcdl, where many unchanged ELS can be observed.
Table 2: The mean energy transitions and ELS time of 54 cases.
@@
Number ascended ELS 7.2 7.8 +0.6 (8.9%)
Number descended ELS 63.5 10.6 -52.9 (-83.4%)
Number unchanged ELS 0.3 52.6 +52.3 (158.5 times)
ELS time (sec) $\downarrow$ 84.3 51.9 -32.4 (-38.4%)
@@ 3.4 Model Efficiency
The mean inference time of SPArcdl for generating ELs pre-selection during SPArc plan optimization is $0.05$ seconds on a CPU, representing a significant improvement in speed compared to the performance of the previous method [17], which is $16$ seconds ($p<0.01$). This enhanced efficiency is primarily due to SPArcdl’s capability to simultaneously model the spot-organ and spot-target information across all gantry angles and produce EL pre-selection in a single forward pass.
@@ 4 Discussion
In this study, we proposed SPArcdl, a deep learning method to rapidly pre-select EL that maintain high treatment quality and significantly reduce ELS time in PAT, without relying on supervised training data. The proposed method is tested on $54$ NPC patient cases and evaluated against a published method SPArcps.
The data preprocessing involves matrices creation of $M_{CTV}$ and $M_{CTV-BRS}$, in which $M_{CTV}$ represents the number of proton spot intersecting the CTV and $M_{CTV-BRS}$ indicates the spots intersecting CTV but not the brainstem. The values of both both matrices are structured by gantry angles and EL. SPArcdl takes in these matrices and WET vectors to generate the optimal EL of all gantry angles. A tri-objective loss function is designed to encourage the number of spots on CTV, penalize the spots on brainstem and discourage time-consuming ELS. The plan quality of SPArcdl is comparable or better than that of SPArcps, and ELS time has been largely improved.
@@
One limitation of SPArcdl is the instability of its predictions, which is largely attributed to the complexity involved in tuning the loss weight parameters. Future work will focus on developing an automated approach to replace the current manual and time-consuming process of objective weight tuning. Another limitation of this study stems from computational resource constraints on the research platform, which necessitated sampling the full-arc gantry angles at 5-degree intervals. We plan to enhance computational capacity to enable the analysis of a greater number of gantry angles in future investigations. Additionally, the current approach only incorporates spot counts for the clinical target volume (CTV) and brainstem, along with the mean water equivalent thickness (WET) value for each angle. Other potentially informative features—such as beamlet-level characteristics, geometric parameters, and a broader range of WET-related features [15]—were not considered. Incorporating these features as model inputs may lead to improved performance.
@@ 5 Conclusion
The proposed SPArcdl is able to achieve comparable or superior plan quality for delivery efficient proton arc therapy plan optimization, compared to the previous work SPArcps. The proposed spot-count representation establishes a foundation for unsupervised deep learning application in EL pre-selection of PAT plan optimization. We hope this work can inspire more deep learning-based methods in this field.

### v2

@@ 2.4 Evaluation
A total of $35$ patients who underwent radiotherapy for NPC were retrospectively selected for evaluation by SPArcps and SPArcdl. The performance of SPArcdl across all five validation datasets ($35$ cases) were aggregated for assessment. We calculate metrics mean for quantitative evaluation and select four representative cases for qualitative evaluation. Two structures, CTV and brainstem, are selected for evaluation. Full arc gantry angle selection was employed, with an angular sampling interval of $5$ degrees due to the constraints of computing power. Our energy selection method depended on matRad supplying the initial complete set of EL that span the entire treatment target longitudinally for every gantry angle. A spot spacing of $5$ mm was used in both the beam and lateral directions. Dose calculations were performed with a resolution of $3\times 3\times 3$ $mm^{3}$, and the voxel size of the CT images was $1.36\times 1.36\times 3$ $cm^{3}$. The prescription dose is $54$ Gy/CTV delivered in $30$ fractions, making $1.8$ Gy per fraction.
@@ 2.4.1 Plan quality evaluation
The performance of SPArcdl in respect to plan quality and robustness of the PAT plan generated by the predicted EL is evaluated. The mean value of D98, D2, homogeneity index (HI), conformity index (CI), mean and max dose of brainstem are calculated for comparison between SPArcdl and SPArcps.
HI quantifies how uniformly the prescribed radiation dose is distributed within the target volume. A low HI value, ideally close to 0, indicating a more homogeneous dose distribution, suggesting less dose variation within the target. Eq.(5) presents the formula of HI described in [27].
$HI=\frac{D_{2}-D_{98}}{D_{prescription}}*100,$ (5)
@@ 2.4.2 EL switch time and model efficiency analysis
A commonly emphasized objective in existing studies is the reduction of treatment delivery time [28, 17, 12, 15, 30]. The duration of arc delivery is influenced by various machine-specific factors, such as the number and order of EL, spot positions, and gantry angles. Among these, ELS time often contributes the most to overall delivery time, especially when the number of energy layers and energy switch-ups is not minimized [28, 17, 12]. The ELS time is evaluated and analyzed for both SPArcps and SPArcdl. The efficiency to generate EL pre-selection, or the inference time of SPArcdl, is also reported.
2.4.3 Robustness evaluation
The robustness of the generated plan is assessed by dose volume histograms (DVHs) with bands and Root-Mean-Square Dose-Volume Histograms (RVH). They effectively capture the sensitivity of dose distributions to uncertainties such as patient setup errors and proton range variations. DVH bands illustrate the variability in dose coverage across multiple uncertainty scenarios, where narrower bands indicate greater dose stability or a more robust plan. RVH complements this by quantifying dose variation using the root-mean-square deviation across scenarios. A smaller RVH AUC reflects a more consistent and resilient dose distribution under uncertainty.
Table 1: The mean dosimetric performance comparison of 35 cases of SPArcps and SPArcdl; D98: the minimum doses received by $98\%$ of the CTV; D2: the minimum doses received by $2\%$ of the CTV; CI: CTV conformity; HI: CTV homogeneity; Mean and Max: the mean dose and max dose of brainstem.
Structure Type SPArcps SPArcdl (Ours)
[…]
Figure 2: Comparison of various metrics performance of $35$ cases between SPArcdl and SPArcps plans. Each diagram shows the relationship between GTV volume and the respective metric, displaying kernel density estimate plots and scatter points for both methods.
@@ 3.1 Plan quality
Table 1 demonstrates the quantitative dosimetric performance comparison of SPArcdl and SPArcps, including D98, D2, CI, HI, mean and max dose of brainstem. With regard to target coverage, the CI of SPArcdl is 0.70, higher than that of the SPArcps ($p<0.01$), suggesting SPArcdl achieves better dose conformity to the target. Compared to SPArcps, the D98 value of SPArcdl increased from $1.46$ to $1.72$ ($p<0.01$) and D2 decreases from $1.99$ to $1.86$ ($p<0.01$). SPArcdl plan presents a significantly ($p<0.01$) lower average HI compared to the SPArcps, indicates more homogeneous dose distribution of SPArcdl. With regard to OAR sparing, the max and mean dose of brainstem (BRS) is reported in Table 1. Both metrics of BRS are lower in SPArcdl plan, with a difference of $0.08$ ($p<0.01$) in max BRS dose and $0.25$ ($p<0.01$) in mean BRS dose, suggeting improved sparing and better protection of BRS by SPArcdl.
@@
A more intuitive performance comparison of SPArcdl (blue dots) and SPArcps (orange triangles) plans is demonstrated in Figure 2, in which the distribution of CI, HI, D98, D2, mean and max dose of brainstem are presented against GTV using scatter plots and kernel density estimate plots. In plot (a), the HI values of SPArcps plan range from about $0.4$ to $2$, with a dense cluster around from $0.4$ to $1$, while those of SPArcdl plan are lower and with a tighter cluster, mostly between $0.2$ and $0.4$, suggesting improved homogeneity of the proposed method. Also in plot (d), the D98 values of SPArcdl plan are higher, mostly between $1.6$ and $1.8$, with a dense cluster around $1.7$ to $1.8$, indicating better minimum dose coverage. The CI values of SPArcdl are higher in plot (b), mostly between $0.6$ and $0.8$ with a dense cluster around $0.7$, while those of SPArcps ranges from $0.5$ to $0.7$, showing better conformity of SPArcdl to the target volume. In plot (e) and (f), both MeanBRS and MaxBRS of SPArcdl exhibits a lower range, suggesting better protection of brainstem. In particular, the MeanBRS values of both methods increase with the GTV, suggesting it is more difficult to spare OAR while GTV is large, which is consistent with expectations. The observations and findings of Figure 2 align with the insights of Table 1. Three cases are selected for qualitative evaluation in Figure 3(b) which exhibit comparable CI and HI performance of SPArcdl and SPArcps. Yet, the brainstem mean dose of SPArcdl is consistently lower than that of SPArcps.
Figure 3: Comparison of SPArcdl and SPArcps plans. (a) The predicted EL delivery sequences of SPArcps and SPArcdl. (b) The dose map overlaid on CT slice with six metrics for each case in nominal distribution. (c) The DVH band and RVH curve of the corresponding cases for robustness analysis. BRS stands for brainstem.
The relatively denser cluster generated by SPArcdl in the six plots of Figure 2 as well as the lower stardard deviation of SPArcdl’s performance arocss the six metrics in Table 1 indicates greater consistency of SPArcdl’s performance. This suggests that SPArcdl not only improves the average dosimetric performance but also ensures more consistent and reproducible results in EL pre-selection.
3.2 EL switch time and delivery sequence
Figure 3(a) demonstrates the prediction of EL delivery sequence of three cases by SPArcdl and SPArcps. The predicted EL of SPArcdl are in the range between $90$ and $140$ Mev, while those of SPArcps spans from approximately $110$ to $160$ Mev, suggesting a more conservative energy selection strategy by SPArcdl. The steeper diagonal trangle lines in the plots indicate SPArcps utilizes a broader spectrum of energy layers, while many data points by SPArcdl cluster in horizontal segments, reflecting unchanged energy layers over multiple gantry angles.
@@ 3.2 EL switch time and delivery sequence
The performance of the model is also evaluated in terms of mean ELS time in Table 2. While many previous works have found using descended ELS is more efficient than ascended ELS [17, 12, 25], the results of SPArcdl unintentionally reveals using unchanged ELS may be more time-wise efficient than descended ELS. Unchanged ELS refers to the scenario in which the EL remains unchanged when the gantry angle transitions to the next position. In such cases, the ELS time is effectively zero, which is more efficient than that required by both ELS upward or downward transitions. Table 2 presents the mean ELS associated with ascended, descended and unchanged transition of SPArcdl and SPArcps. While the mean ascended ELS of SPArcdl is $7.7$, slightly higher than that of SPArcps ($6.6$), the mean unchanged ELS of SPArcdl is $53$, which is dramatically higher than that of SPArcps or $0.3$. Compared to SPArcps, the number of descended ELS of SPArcdl has vastly shifted to the unchanged ELS. The notable advantage of SPArcdl lies in its higher frequency of unchanged ELS. This advantage not only offsets the the cost of slightly higher EL in ascended transitions but also contributes to a gain of energy switch time of $30.1$ seconds, an $37.2\%$ reduction from that of SPArcps. Figure 3(a) demonstrates EL generated by SPArcdl of three cases, where many unchanged ELS can be observed.
Table 2: The mean energy transitions and ELS time of 35 cases.
@@
Number ascended ELS 6.6 7.7 +1.1 (16.7%)
Number descended ELS 64.1 10.3 -53.8 (-83.9%)
Number unchanged ELS 0.3 53.0 +52.3 (176.7 times)
ELS time (sec) $\downarrow$ 81.1 $\pm$ 7.7 51.0 $\pm$ 11.7 -30.1 (-37.2%)
3.3 Robustness
Comparative analysis of the DVH bands and RVH curves between SPArcdl and SPArcps plans for two representative cases is presented in Figure 3(c). While the DVH bandwidths for both CTV and brainstem demonstrate comparable characteristics between the two methods, notable differences emerge in their dose distribution patterns. SPArcdl exhibits higher-than-prescription doses in certain CTV regions, whereas SPArcps maintains doses within prescribed limits. This disparity suggests increased dose coverage variability under uncertainties for SPArcdl, indicating its reduced plan robustness compared to SPArcps. Analysis of RVH curves reveals that SPArcdl generates comparable area under curve (AUC) for brainstem, but larger area for CTV, further substantiating increased dose variation under uncertainty. The inferior robustness performance of SPArcdl can be attributed to the absence of robustness-specific objectives in its optimization framework, consequently limiting the network’s capacity to improve robustness-related metrics such as DVH bands and RVH AUC.
@@ 3.4 Model Efficiency
The mean inference time of SPArcdl for generating EL pre-selection for SPArc plan optimization is $0.05$ second on a Intel Core i7 CPU, representing a significant improvement in speed compared to the performance of SPArcps, which is $13.5$ seconds ($p<0.01$) on the same CPU. This enhanced efficiency is primarily due to SPArcdl’s capability to simultaneously model the spot-organ and spot-target information across all gantry angles and produce EL pre-selection in a single forward pass.
@@ 4 Discussion
In this study, we proposed SPArcdl, a deep learning method to rapidly pre-select EL that maintain high treatment quality and significantly reduce ELS time in PAT, without relying on ground truth required by supervised approach. The proposed method is tested on $35$ NPC patient cases and evaluated against a published method SPArcps.
A novel data representation method, spot-count representation, involves the creation of matrices $M_{CTV}$ and $M_{CTV-BRS}$. Specifically, $M_{CTV}$ represents the number of proton spots intersecting the CTV, while $M_{CTV-BRS}$ indicates the spots intersecting CTV but avoiding the brainstem. Both matrices are structured by gantry angles and energy layers and serve as proxies for radiation dose coverage. SPArcdl processes these matrices along with WET vectors to determine the optimal energy layer for each gantry angle. The model employs a tri-objective loss function that maximizes target coverage by encouraging spot placement on CTV, minimizes organ-at-risk exposure by penalizing spots on the brainstem, and improves delivery efficiency by discouraging time-consuming energy layer switches. In terms of nominal dose distribution, SPArcdl achieves comparable or superior plan quality in CTV coverage and brainstem sparing compared to SPArcps, with significantly reduced ELS time. However, further improvements in SPArcdl’s plan robustness are needed.
@@
While SPArcdl demonstrates promising results, its limitations merit careful consideration. First, the model currently lacks explicit mechanisms to address treatment uncertainties, particularly those arising from patient setup variations and intrafractional motion, which may impact the robustness of the selected EL in clinical settings. Second, computational resource limitations necessitate the sampling of full-arc gantry angles at 5-degree intervals instead of the more precise 2-degree intervals. Future research will focus on integrating robustness-aware strategies to improve both the accuracy and applicability of SPArcdl in real-world radiotherapy planning scenarios.
@@ 5 Conclusion
In this study, we have demonstrated that SPArcdl achieves comparable or superior plan quality in nominal distribution while significantly improving delivery efficiency and accelerating inference speed in EL pre-selection task compared to the established SPArcps method. The novel spot-count representation introduced here provides a strong foundation for implementing unsupervised deep learning methods in this task. An interesting finding is the experimental results suggest a high frequency of unchanged ELS deployments may be more efficient than a high frequency of descended ELS. Even though SPArcdl has limitation regarding plan robustness, this research represents the first step towards integration of deep learning techniques in EL pre-selection for PAT plan optimization and may catalyze future developments in deep learning-based PAT plan optimization strategies.
Data availability statement
The data cannot be made publicly available upon publication because no suitable repository exists for hosting data in this field of study. The data that support the findings of this study are available upon reasonable request from the authors.
Ethical statement
This study was approved by the Institutional Review Board at the Tongji Medical College of Huazhong University of Science and Technology (ethics number: [2020] (0361)). All methods were performed in accordance with the relevant guidelines and regulations. Written informed consent was obtained from all authors and participants.